The Spontaneous Galvanic cell Tin | Tin ion (0.024M) ‖ Tin ion (0.064M) | Tin develops an emf of 0.0126 V at 250C, Calculate the Valency of Tin.
1
Expert's answer
2019-09-12T03:31:45-0400

"E_{cell}=\frac{0.0591}{n}log\frac{C_2}{C_1}"

"0.0126=\frac{0.0591}{n}log\frac{0.064}{0.024}"

"n=1.998 \approx 2"

Hence the valency of "Ti" is 2.